Please help me find the first three terms in this expansion:
(x+3y)^20

Answer by stanbon(75887)

find the first three terms in this expansion:
(x+3y)^20 = x^20 + 20C19*x^19(3y) + 20C18*x^18(3y)^2 +....
= x^20 + 20*3 x^19y + 180*9 x^18y^2.....
---
= x^20 + 60x^19h + 1620x^18y^2....
========================================
Cheers,
Stan H.
=============